# Usable XAI: 10 Strategies Towards Exploiting Explainability in the LLM Era

### v1

@@ 2.1.1 Attributing Inputs for Label Classification
Table 1: Time complexity analysis on different attribution methods for the generative task.
@@
To summarize, the traditional explanation methods are not always suitable for LLMs. In particular, the perturbation-based and gradient-based explanations are relatively easy to extend for attributing LLM responses to the input prompts, while the surrogate-based and decomposition-based methods become significantly challenging to do so. Specifically, surrogate-based methods suppose that an explainable small model could approximate the decision boundary of the target model around a local example, but there are limited explainable models for the text generation task. Meanwhile, decomposition-based methods require designing decomposition strategies for different layers, which is challenging for big LLM architectures. Another primary concern is their significant demand for computing resources. Given an $N$-words input prompt and an $M$-words output response, we present the time complexity of several representative explanation methods in Table 1. It demonstrates that existing methods either require a large number of forward operations or backward operations. Therefore, improving the efficiency of the attribution-based explanation is an important direction for future research and development.
2.1.2 Attributing Inputs for Text Generation
Figure 2: An example of attribution saliency map between input and output tokens.
The explanation of generative models can be defined as attributing the overall confidence $p(\hat{y}|x)$ to the input $x$, where $\hat{y}$ denotes the generated response $\hat{y}=[\hat{y}_{1},...,\hat{y}_{M}]$ with $M$ words. One method to achieve this is by treating the text generation process as a sequence of word-level classification tasks. This perspective allows for the application of existing classification-based explanation techniques to assess the influence of each input word $x_{n}$ in relation to each output word $\hat{y}_{m}$, resulting in a corresponding attribution score $a_{n,m}$. After gathering the attributions $a_{n,m}$ for $m=1,...,M$, we perform an aggregation to determine the overall contribution of each input word $x_{n}$. This is accomplished by aggregating the individual attributions for all output words corresponding to the input word, denoted as $a_{n}=\mathrm{Aggregate}([a_{n,1},...,a_{n,M}])$. The simplest approach for this aggregation is to average the attributions assigned to each input word across the different output words (Selvaraju et al., 2016). However, Wu et al. (2023) observe that attribution scores from different output words are not inherently comparable. For example, the attribution scores for function words (e.g., “the”, “is”, “have”) are often disproportionately larger than the scores for content words with clear semantic meaning (e.g., verbs and nouns). Therefore, it is necessary to normalize the scores prior to the aggregation, so that the scores $[a_{n,1},...,a_{n,M}]$ become comparable for $1\leq m\leq M$. Figure 2 plots the normalized scores of an example case, where each index in the Y-axis refers to an input prompt token, while that in the X-axis is an output response token. A greater normalized attribution score is brighter. In this example, the user attempts to direct the model to output information that does not exist, namely the French president in 1250. The model successfully realizes that this thing does not exist and refuses to answer. The model response can be realized as three parts, “There was no”, “president in France”, and “in 1250”. According to the figure, the first span is generated heavily because of the tokens “Who” and “president”, while the model uses both “France” and “1250” to respond to the second span “president in France”. Finally, the model emphasizes the date “1250” again by referencing the same information from the prompt. Overall, these explanations align with human understanding and highlight the usage of this method in the future. However, current research on attribution-based explaining for generative LLMs is still in its early stages, and only a limited number of methods have been proposed.
2.2 Case Studies: Usability of Attribution Methods for LLMs
Figure 3: A general pipeline of model diagnosis with attribution explanations.
The attribution map offers a partial insight into the operational mechanics of LLMs (Chandrasekaran et al., 2018; Hase & Bansal, 2020; Ye & Durrett, 2022a). Accordingly, we propose a general pipeline that leverages attribution scores to analyze LLM behaviors, as shown in Figure 3. First, given the target LLM and an input prompt, we compute attribution scores of input tokens relative to the output tokens. Second, we extract a feature vector from the attribution map, tailored to the requirements of the diagnostic task at hand. Third, we train a light-weight predictor (e.g., a classifier) to diagnose whether the model behaves appropriately based on the feature vector. In the following, we provide case studies to illustrate how attribution scores could be utilized to assess LLM response quality (Adlakha et al., 2023).
2.2.1 LLM Response Quality Evaluation with Explanations
This case study explores the use of attribution-based explanations as evidence for assessing the quality of LLM-generated responses. Here, “quality” is specifically measured by the responses’ accuracy. We hypothesize that responses generated from correct rationales are likely to be more accurate. Our method involves comparing the model’s explanations with underlying rationales as a means to evaluate response accuracy. This approach could enhance the reliability of conversational systems, which is crucial for applications such as medical question answering.
Dataset.
We employ the Multi-Sentence Reading Comprehension (MultiRC) dataset (Khashabi et al., 2018) in this case study. MultiRC presents a more significant challenge than other datasets: it requires the system to answer questions based on multiple sentences from a given paragraph. The answers may not directly correspond to specific sentence spans, reflecting a more realistic scenario of modern LLMs. Each MultiRC instance is associated with a human-annotated rationale indicating which sentences are relevant for answering the question. Our study specifically employs its test subset with 950 samples in total. We consider 80% in which are the training set and we report the results evaluated on the rest samples.
Implementation.
We apply the Importance Density Score method (Wu et al., 2023) to estimate the importance of each word in the provided paragraph. Specifically, given $N$-word paragraph $X$ and the $M$-word response $\hat{Y}$, the attribution from each input word $x_{n}$ to each output word $\hat{y}_{m}$ is defined as $a_{n,m}=\frac{\partial f(y_{m}|Z_{m})}{\partial\,\mathbf{E}_{i}[x_{n}]}\cdot\mathbf{E}_{i}[x_{n}]^{\top}$, where $Z_{m}$ indicates the formatted input prompt concatenated with the first $m-1$ response word, $f$ is the language model, and $\mathbf{E}_{i}[x_{n}]$ indicates the input static embedding of $x_{n}$. This pairwise attribution score is then normalized as $\widetilde{a}_{n,m}=\lceil L\times a_{n,m}/\text{max}_{n^{\prime}}(a_{n^{\prime},m})\rceil$. Any normalized attribution $\widetilde{a}_{n,m}$ is equal or less than $b$ will is forced to 0. The overall attribution score on word $x_{n}$ is defined as its attribution density, i.e., $a_{n}=||[a_{n,1},...,a_{n,M}||_{1}/||[a_{n,1},...,a_{n,M}]||_{p}$. In our experiments, we let $L=10$, $b=2$, $p=5$, and consider Vicuna-7B-v1.1 as our language model $f$. By averaging these word attribution scores, we reach the importance of each sentence. The top-K sentences with the highest importance scores are selected as the explanation for each instance. The explanation is then concatenated with the output response and fed to a classifier to predict whether the output is correct or not. We initialize the classifier with DistillBERT-base (Sanh et al., 2019), and train it over 3 epochs with learning rate $5e^{-5}$ and weight decay $1e^{-3}$. To compare with this, we also consider the Full Paragraph or the human-annotated rationales as the explanation part to train the classifier. Following previous studies (Khashabi et al., 2018; DeYoung et al., 2019), we evaluate the accuracy of the generated responses by exactly matching all correct answers with the generated responses. Table 2 reports the macro precision, recall, F1, and AUC scores on the test set over five random seeds.
Results.
Table 2: Utilizing attribution for response quality evaluation
Setting Precision Recall F1 AUC
Random 49.40 51.79 49.61 49.03
Human Rationale 68.73 66.88 67.57 73.11
Full Paragraph 58.02 58.47 56.89 63.44
Attribution (ours) 63.25 67.69 64.12 71.53
In Table 2, we observe that isolating the rationales from the full context could best help the classifier identify the response quality. It is evident that explanations align more closely with human-annotated rationales when associated with correct responses. In particular, the precision and recall metrics for explanations derived from correct answers surpass those associated with incorrect responses. This finding serves as empirical evidence underscoring the effectiveness of attribution-based explanations in estimating the accuracy of responses generated by LLMs.
2.2.2 Hallucination Detection with Attribution Explanations
This case study explores a different aspect of LLM generation quality, focusing on the presence of hallucinations in LLM-generated responses. We show that attribution-based explanations can serve as indicators to detect LLM hallucinations. Hallucinations are defined as responses that contain information conflicting with or unverifiable by factual knowledge (Li et al., 2023c; Ji et al., 2023). For instance, if a model is asked about a fictitious entity like “King Renoit” and responds with a narrative about “The Three Musketeers”, claiming it pertains to the nonexistent king, it illustrates a hallucination. This tendency, particularly pronounced in instruction-tuned models, stems from their earnest endeavor to fulfill user requests. The issue often arises when direct commands (“tell me a story”) significantly influence the generation process, while the instruction’s subject (“about King Renoit”) is neglected. Based on this insight, we develop a hallucination detector according to the distribution of attribution scores over different types of prompting words.
Dataset.
We use the Hallucination Evaluation Benchmark (Li et al., 2024a) in this case study. Each instance from this dataset includes an input prompt, a response generated by ChatGPT (Ouyang et al., 2022), and the knowledge related to the response. Each piece of knowledge has a human annotation about whether the knowledge is valid or not. Each instance having at least one invalid knowledge is considered a hallucination response. This study focuses on 632 less confused examples from the benchmark, each of which has almost all knowledge correct or almost all wrong. We randomly select 80% of the samples to form a training set and the rest form the testing set.
Implementation.
Given a query prompt and its ChatGPT response, we aim to build a classifier to detect if the response contains hallucination. Since the gradients of ChatGPT is inaccessible, we apply the Vicuna-7B model as the substitute to compute the attribution scores. Specifically, we adopt Importance Density Scores (Wu et al., 2023) to compute the attribution scores of Vicuna. We then use the NLTK package to identify the part-of-speech (POS) tag of each query word. Finally, each query-response pair is represented with an 82-dimensional vector, where each dimension indicates the average attribution scores over a certain type of POS tagging. We develop a Support Vector Machine classifier based on the POS tagging attribution scores on the training set and report the precision, recall, F1 scores on the hallucination samples, as well as accuracy over all samples in Table  3. To compare with this, we also leverage a fine-tuned model and a prompting-based method to serve as the hallucination detector baselines.
Results.
Table 3: Utilizing attribution for hallucination detection on ChatGPT responses.
Method Language Model Precision Recall F1 Accuracy
Random - 88.41 50.34 64.11 50.59
FacTool GPT-4 95.30 72.93 82.62 73.04
Vectara DeBERTa-base 90.29 60.54 72.40 59.45
AttrScore (ours) Vicuna-7B 90.15 74.21 81.36 70.20
Mistral-7B 88.74 75.04 81.26 69.57
In the table, we first observe that all methods have demonstrated a greater performance than the Random strategy, indicating that the proposed methods and the baselines are valid hallucination detectors. In addition, we observe that the proposed methods achieve a competitive performance with FacTool using GPT-4, establishing a new benchmark in this domain. It is worth noting that, instead of using GPT-4, the attribution score for our method is calculated using a smaller 7-billion-parameter language model. This demonstrates the practicality and efficiency of our approach in weak-to-strong generalization as we could diagnose large language models with smaller models. Future work could consider extracting more effective features and using more powerful classifiers.
2.3 Challenges
2.3.1 How to Identify and Explain the Semantics of Output?
The attribution function $a_{n}=p(\hat{y}|x)-p(\hat{y}|\widetilde{x}_{n})$ is no longer faithfully attributing the model predictions in the human-interested semantic level since the model could express the same semantic meaning with various responses. Specifically, the model could assign a lower confidence to its original response than the new one, while both responses share the same semantic meaning. This is a significant difference compared with the traditional classification problem, where target label sets are manually designed so that a lower $p(\hat{y}|\cdot)$ indicates the model is less confident in predicting a specific semantic concept. Taking the sentiment analysis task as an example, an LLM may generate two different responses sharing the same predicted concepts, such as “it is a positive review” and “the audience thinks the movie is great”. Current attribution-based explanations concentrate on the literal changes in generated responses, but they do not study how the semantic meanings of these responses change. Therefore, they do not provide sufficient explanations of model-generated responses at a semantic level. In this case, the semantic level is which words of the input review lead the model to believe it is positive. Future work may tackle this challenge by proposing metrics to evaluate the semantic differences in responses.
Figure 4: Review of interpretation methods for LLM components and their applications. We categorize methods according to the target LLM modules: self-attention layers and feed-forward layers.
2.3.2 Explaining LLM Predictions Beyond Attribution
The versatility of LLM generation will inspire a variety of explanation paradigms beyond traditional attribution methods. Attribution methods aim to explain model output with the contribution of input features. This explanation task is meaningful for conventional machine learning (ML) models whose outputs are usually individual decisions with clear formats (e.g., classification, regression, object detection). The decisions are highly dependent on the input features. However, LLMs differ from traditional ML models in two aspects. First, the generative process of LLMs is stochastic, e.g., we may obtain different outputs by running an LLM twice even with the same input. Second, LLMs encode rich knowledge within their parameters, which are independent of inputs. These unique properties give rise to novel explanation paradigms. For the first aspect, an explanation task of interest would be to understand the uncertainty of LLM generation. For example, researchers (Ahdritz et al., 2024; Varshney et al., 2023; Su et al., 2024) leverage the prediction perplexity to check whether the LLM is confident during generation, identifying potential errors in less confident predictions. Second, attributing LLM predictions to their encoded knowledge instead of input patterns could provide a new perspective. Some researchers (Yin et al., 2024a) propose the knowledge-boundary detection task to detect whether a model has specific knowledge for a given question. Nevertheless, it did not attribute the prediction to specific knowledge, so humans cannot verify the prediction process with their results yet.
3 LLM Diagnosis and Enhancement via Interpreting Model Components

### v2

@@ 2.1.2. Attribution Methods for Generation
Table 1. Time complexity analysis on different attribution methods for the generative task.
@@
Another category of attribution approaches leverages the generative capabilities of LLMs to provide citations to one or more text passages that support the output they generate. The attribution can be obtained by explicitly prompting LLMs to quote the curated sources of information (Weller et al., 2024). The cited contents can originate either from external knowledge sources or from the LLMs’ own memory. For instance, Sun et al. (2023) introduces RECITE, which first retrieves one or more relevant passages from the LLMs’ memory via sampling and then generates the final response based on these passages. Conversely, many systems assume the presence of an external retrieval corpus. In these cases, the LLM output includes both an answer string and a reference to a short segment of text from the corpus that substantiates the answer (Bohnet et al., 2022; Gao et al., 2023c; Asai et al., 2023a). Nevertheless, LLMs may hallucinate and produce fabricated information. To address this, various benchmarks have been developed to assess the faithfulness of LLM-generated attributions, i.e., whether the generated statements are fully supported by the cited references. For details on evaluation methods, we refer readers to Section 11.
2.2. Usability of Attribution Methods for LLMs
Explanations provide a lens to evaluate whether a model’s predictions are grounded in reasonable rationale and to identify opportunities for improving prompt instructions. Based on this, we discuss two scenarios where LLMs can benefit from attributions: model diagnosis and model enhancement.
  • Model Diagnosis. By comparing the LLM’s output with attribution results, it is possible to detect ungrounded or unsupported information. Formally, let $x$ denote the input, $\hat{y}=f(x)$ be the target model’s output, and $a$ be the attribution. The goal is to construct a detector $g_{\mathrm{detect}}(a,x,\hat{y})\in\{-1,+1\}$ where $+1$ indicates that the prediction can be trusted, while $-1$ means the opposite. Here $g_{\mathrm{detect}}$ can be implemented as an LLM. For instance, Gao et al. (2023a) proposes a framework called RARR, which leverages attribution to identify unsupported or misleading content generated by LLMs. In particular, the attribution step is automated by using a query generator to produce comprehensive questions about different aspects of the text that need verification. In addition, Agrawal et al. (2024) proposes ”consistency checks”, which asks the LLM to perform attribution multiple times for the same output. Inconsistencies in the attribution results serve as indicators of potential issues with the factuality of the LLM’s output.
  • Model Enhancement. Attribution results can serve as additional information to refine prompts, thereby improving LLM performance. For instance, Sun et al. (2023) proposes ”recitation” as a new two-step approach where an LLM $f$ first explicitly recalls relevant factual knowledge $a$ from its own memory before generating an answer for input $x$. The LLM inference is augmented as $\hat{y}=f(x,a)$ instead of the standard $f(x)$. This recitation step allows the model to better access and verify its stored knowledge, similar to how a student might first recite relevant facts they’ve learned before answering an exam question, leading to more accurate responses. In addition, Krishna et al. (2024) proposes a framework called AMPLIFY, which uses attribution results to improve LLM performance by crafting effective prompts. It selects samples in the validation set that were misclassified by the LLM, and uses another model (e.g., GPT-2 or BERT) to compute their attribution scores with respect to the ground truth labels. These samples and their corresponding explanations are used to construct prompts for future test samples.
Figure 2. Left: An example of attribution saliency maps. Right: The pipeline of making use of attribution results to improve LLMs.
2.3. Challenges
We discuss two major challenges regarding attribution methods for explaining LLMs, including efficiency and faithfulness. First, according to our analysis in Table 1, many traditional attribution methods are computationally expensive when adapted to LLMs. Furthermore, these methods were primarily developed for tasks in the context of classification or regression, while the more urgent need lies in designing attribution methods tailored for LLM generation. The diverse and dynamic nature of LLM generation demands the development of novel explanation paradigms that go beyond existing frameworks. Second, while some approaches utilize the generative capabilities of LLMs to produce free-text attributions, the faithfulness of these attributions remains underexplored. Ensuring that the generated explanations truly reflect the model’s reasoning process is a challenge to be tackled for building trust and reliability in LLM explanations.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, rectangle, draw, font=, rounded corners, align=center, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=2pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, text centered, , [Interpreting Model Internals,ver,fill=gray!20,text centered,text width=10em [Explanation of
Model Components,ver,fill=orange!40,text centered,text width=7em [Importance Analysis on
Single Component,fill=orange!10,text centered,text width=6em [Contextual Impacts: (Geva et al., 2022), (Ferrando et al., 2023), (Wang et al., [n. d.]), (Elhage et al., 2021), (Meng et al., 2022a), (Wang et al., 2022a), (Mohebbi et al., 2023)., leaf2, align=left] [Global Impacts: (Geva et al., 2021), (Dar et al., 2023), (Wu et al., 2023b)., leaf2, align=left] ] [Circuit Analysis on
Multiple Components,fill=orange!10,text centered,text width=6em [(Cammarata et al., 2020), (Wang et al., 2022a), (Hanna et al., 2023), (Geiger et al., 2021), (Conmy et al., 2023) (Sarti et al., 2023), (Hanna et al., [n. d.]), (Marks et al., 2024)., leaf2, align=left] ] [Usability,fill=orange!10,text centered,text width=6em [Knowledge Editing: (Dai et al., 2022), (Wang et al., 2022b), (Elhage et al., 2021), (Meng et al., 2022a), (Meng et al., 2022b), (Hase et al., 2024), (Li et al., 2024a),
(Gupta et al., 2024), (Wu et al., 2023a), (Wang et al., 2024c)., leaf2, align=left] [Model Pruning: (Dalvi et al., 2020), (Men et al., 2024), (Gromov et al., 2025), (Dalvi et al., 2020), (Fan et al., 2024)., leaf2, align=left] ] ] [Explanation of
Latent Representations,ver,fill=orange!40,text centered,text width=7em [Non-Training
based Methods,fill=orange!10,text centered,text width=6em [Projecting $\mathbf{x}^{l}$ to Output Embeddings: (nostalgebraist, 2020), (Jastrzebski et al., 2018), (Belrose et al., 2023), (Din et al., 2024)., leaf2, align=left] [Feeding $\mathbf{x}^{l}$ as Input Embeddings: (Chen et al., 2024b), (Dumas et al., 2024), (Pan et al., 2024)., leaf2, align=left] ] [Supervised Training
based Methods,fill=orange!10,text centered,text width=6em [Identifying Concept Existence: (Hewitt and Manning, 2019),(Chen et al., 2021), (Liu et al., 2019), (Hewitt and Manning, 2019), (Burns et al., 2022),
(Zou et al., 2023a), (MacDiarmid et al., 2024),(Zheng et al., 2024), (Hewitt and Liang, 2019), (Pimentel et al., 2020), (Voita and Titov, 2020)., leaf2, align=left] [Measuring Concept Importance: (Kim et al., 2018), (Wei et al., 2021), (Bai et al., 2022), (Nicolson et al., 2024)., leaf2, align=left] ] [Unsupervised Training
based Methods,fill=orange!10,text centered,text width=6em [Orthogonal Assumption on Concepts: (Millidge and Black, 2022), (Wu et al., 2023b)., leaf2, align=left] [Sparse Assumption on Concepts: (Olshausen and Field, 1997), (Cunningham et al., 2023), (Bricken et al., 2023)., leaf2, align=left] ] [Usability,fill=orange!10,text centered,text width=6em [Data Pre-processing: (Ge et al., 2024), (Xu et al., 2023c), (Li et al., 2024c), (Yang et al., 2025b)., leaf2, align=left] [Training-Time Regularization: (Yin et al., 2024a), (Wu et al., 2025b)., leaf2, align=left] [Inference-Time Control: (Wu et al., 2025c), (Huang, 2024), (Stolfo et al., 2024), (Rimsky et al., 2024), (Arditi et al., 2025), (Bhattacharjee et al., 2024),
(Zou et al., 2023a), (Rimsky et al., 2024), (Chen et al., 2023a), (Ahdritz et al., 2024), (Singh et al., 2024), (Zou et al., 2023a), (Scialanga et al., 2025)., leaf2, align=left] ] ] ]
Figure 3. Taxonomy of explanation methods for LLM internals, including explaining model components and hidden representations.
3. LLM Diagnosis and Enhancement via Interpreting Model Internals